Craylon Corp has two service departments, S1 and S2, and two production departments, P1 and P2.

The data for April were as follows:

Services provided to:

Activity

Costs

S1

S2

P1

P2

S1

$100,000

20%

40%

50%

S2

$80,000

10%

55%

25%

Fixed Costs

P1

$400,000

P2

$500,000

Required:

Allocate the service departments' (S1 & S2) costs to the two operating divisions using the reciprocal method.
Answer #1
To
From S1 S2 P1 P2
Department cost $          1,00,000 $      80,000 $   4,00,000 $      5,00,000
S1 $         -1,11,134 $      20,206 $       40,412 $          50,515
S2 $              11,134 $ -1,00,206 $       61,237 $          27,835
Total Cost $                       -   $                -   $   5,01,649 $      5,78,351
Workings:
Let Total Administrative department cost be x and Accounting Department be y
x = 100000 + (10/90 * y)
y = 80000 + (20/110 * x)
By solving above:
y = 80000 + (20/110 * x)
y = 80000 + (20/110 * (100000 + (10/90 * y)))
y = $ 100206
also,
x = 100000 + (10/90 * y)
x = 100000 + (10/90 * 100206)
x = $111134